OpenStudy (anonymous):

Solve the differential equation with Fourier transformations: y'(t)+t*y(t)=0 ,y(0)=1

OpenStudy (solomonzelman):

y+t•y=0 Find the integrating factor

OpenStudy (solomonzelman):

Anytime, you are dealing (in general) with: \(\large\color{#000000 }{ \displaystyle y'+q'(x)\cdot y=p(x)}\) You are multipling the equation times \(\large\color{#000000 }{ e^{q(x)}}\) all through the equation. \(\large\color{#000000 }{ \displaystyle y'e^{q(x)}+q'(x)e^{q(x)}\cdot y=p(x)e^{q(x)}}\) And the left side would become, \(\large\color{#000000 }{ \displaystyle \frac{dy}{dx}\left[ye^{q(x)}\right]=p(x)e^{q(x)}}\)

OpenStudy (solomonzelman):

And at that point you are capable of integrating both sides.

OpenStudy (anonymous):

Is the integration factor e ^{x ^{2}/2}?

OpenStudy (solomonzelman):

Yes,very good

OpenStudy (solomonzelman):

So multiply \(\large\color{#000000 }{ \displaystyle y'+yt=0}\) times \(\large\color{#000000 }{ \displaystyle e^{t^2/2}}\)

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle y'e^{t^2/2}+ye^{t^2/2}t=0}\)

OpenStudy (solomonzelman):

And then, work the product rule backwards (as I showed).

OpenStudy (anonymous):

Okay I will try!

OpenStudy (solomonzelman):

Sure, go ahead \(: )\)

OpenStudy (solomonzelman):

(if you want to go ahead and integrate both sides, the integral of 0 is equal to C)

OpenStudy (solomonzelman):

(integrate, of course, after you work the product rule backwards.)

OpenStudy (anonymous):

Okay and then I can use Fourier transformations?

OpenStudy (anonymous):

\[e ^{t ^{2}/2}*y=C\]

OpenStudy (solomonzelman):

All you need is (will will consider that general case) BUT, when p(x) is 0. \(\large\color{#000000 }{ \displaystyle y'+w'(x)\cdot y=0}\) And some initial point, lets call it \((4,2)\) You are multipling the equation times \(\large\color{#000000 }{ e^{w(x)}}\) all through the equation. \(\large\color{#000000 }{ \displaystyle y'e^{w(x)}+w'(x)e^{w(x)}\cdot y=0}\) And the left side would become, \(\large\color{#000000 }{ \displaystyle \frac{dy}{dx}\left[ye^{w(x)}\right]=0}\) Integrate both sides, \(\large\color{#000000 }{ \displaystyle ye^{w(x)}=C}\) Solve for y, \(\large\color{#000000 }{ \displaystyle y=\frac{C}{e^{w(x)}}}\) Solve for C, \(\large\color{#000000 }{ \displaystyle 2e^{w(4)}=C}\) ^^ this is constant. So, the answer is (for this abstract problem) \(\large\color{#000000 }{ \displaystyle y=\frac{2e^{w(4)}}{e^{w(x)}}}\)

OpenStudy (anonymous):

I have \[y=\frac{ 1 }{ e ^{t ^{2}/2} }\]

OpenStudy (anonymous):

Thank you! I found it: \[F[y]=e ^{-t^2 /2} \] So the answer is \[y= e^{-w^2/2}*sqrt(2*Pi)\]

OpenStudy (irishboy123):

????

OpenStudy (anonymous):

What don't you understand @IrishBoy123 ?

OpenStudy (irishboy123):

where are the fourier transforms?

OpenStudy (irishboy123):

did i miss that bit?

OpenStudy (anonymous):

Oh wait

OpenStudy (irishboy123):

as in \[\large \hat{f}(\omega ) = \int_{-\infty}^\infty \,dx \qquad f(x)\ e^{- 2\pi i \omega \; x } \]

OpenStudy (anonymous):

\[F[\frac{ 1 }{ \sqrt(2*\pi)} *e^{(-t^2)/2}]=e^{-w^{2}/2}\] So \[F[e^{-t^{2}/2}]=\sqrt{2*\pi}*e^{-w^{2}/2}\]

OpenStudy (anonymous):

@SolomonZelman can you verify my answer?